Use the following paycheck to answer the question.

EMPLOYEE PAY STUB
EMPLOYER NAME EMPLOYEE NAME
Hillsdale Shipping Zachary Smith
EMPLOYER ADDRESS EMPLOYEE ADDRESS
3305 Dupont Square 245 Main Street
Tucson, AZ 72508 Drasco, AZ 72530
EMPLOYER TELEPHONE EMPLOYEE #
870 613 0000 924 2204709
BANK ACCT SSN (LAST FOUR DIGITS)
166008740 1255
PAY PERIOD PAY DATE
07/05/2019 - 07/19/2019 07/26/2019
GROSS EARNINGS RATE HOURS CURRENT TOTAL YEAR TO DATE
REGULAR 18 80 $1440.00 $20,160.00
OVERTIME $0.00 $0.00
OTHER $0.00 $0.00
DEDUCTIONS CURRENT TOTAL YEAR TO DATE
FICA - MEDICARE $20.88 $292.32
FICA - SOCIAL SECURITY $89.28 $1249.92
FEDERAL TAX $172.80 $2419.20
STATE TAX $62.37 $873.18
OTHER $0.00 $0.00
CURRENT GROSS YEAR-TO-DATE GROSS
$1440.00 $20,160.00
CURRENT DEDUCTIONS YEAR-TO-DATE DEDUCTION
$ $
CURRENT NET PAY YEAR-TO-DATE NET PAY
$ $15,325.38


What percent of Zachary's pay do his deductions comprise? Round your answer to the nearest whole percent.
Answer #1

Current

pay= $1440

Deduction=$ 345.33

Net =$ 1094.67

% deduction of gross pay= 345.33/1440*100 =23.98%

≈24%

To date

Pay=$ 20160

Deduction=$ 4834.62

% deduction= 4834.62/20160*100 ≈ 24%